solve system by substitution
-5x-y=-11
-6x+y=-22

Answer by icefreezer7(1)

first you have to solve for the one of the equations.
-5x-y=-11
y=-5x+11
then substitute it in the other equation.
-6x+y=-22
-6x+(-5x+11)=-22 and solve
-11x = -33
x =3
plug in x in any of the equations and you will have x and y.